Weather Modeling: Parallelized and Pedagogical

Skylar Thompson November 20, 2005

Contents

1 Introduction

2 Weather Modeling

• Pre-computers: looking west

• Need more sources here

2.1 Computational 2.2 Parallel Aspects of Weather Modeling

• History back to 1960s on SOLOMONs

• Mesh

• Data type

3 What is Cligen?

• Model vs. Forecast – Short-term vs. long-term – Human vs. computer • Data types

• Compare with MM5

3.1 Parallelization points

• Core statistical equations (marked with PARALLELIZE)

– Calculation of the distance from the mean of precipitation, with units being std.dev.

– Smoothing of monthly rainfall data

– Computation of α, the ratio of half-hourly rainfall to total rainfall

– Computation of χ^{2}, or how well the data fit to a theoretical data

– Computation of Taylor series. I don’t know what it is or what it is used

for, but I could parallelize it.

• Efficiency calculations

• Speedup calculations

4 User interface

• Intuitive

• Easy to make changes

• Fast results

• Explanation?

• Ability to combine multiple stations?

• How fine-grained should the changes be?